OpenStudy (anonymous):

ABCD is a trapezoid with midsegment EF. If AB=16 and EF=34, find CD. A)52 B)18 C)50 D)25

OpenStudy (shadowlegendx):



OpenStudy (shadowlegendx):

Follow this formula. 2Midsegment = Base 1 + Base 2 In this case 2EF = AB + CD

OpenStudy (shadowlegendx):

Now we fill in 2(34) = 16 + CD Now we solve for CD 68 = 16 + CD 52 = CD

OpenStudy (shadowlegendx):

A) 52 is your answer :) @ShadowLegendX me if you need further help, or pm me :P

OpenStudy (shadowlegendx):

get it? @Xpertviper

OpenStudy (anonymous):

Thanks ShadowLegendX

OpenStudy (shadowlegendx):

np :)